OpenStudy (anonymous):

@eSpeX Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = -one halfx + 3 y = -one halfx + 6 y = 2x + 3

OpenStudy (anonymous):

sort of the same thing

OpenStudy (espex):

This one is very much like the one we just did. Yes, the difference between a parallel line and a perpendicular line is that the slope is inverted.

OpenStudy (espex):

So first identify your slope, your \(x_1\) and your \(y_1\)

OpenStudy (anonymous):

like absolute value

OpenStudy (espex):

Like flipping a fraction.

OpenStudy (espex):

So what is the slope of your existing line?

OpenStudy (anonymous):

it is 2

OpenStudy (espex):

Okay, that means then, that your perpendicular line has a slope of \(\frac{1}{2}\)

OpenStudy (anonymous):

ok so up one over 2

OpenStudy (matlee):

it does not turn negative? aw man i put that on my end of course exam today lol

OpenStudy (anonymous):

sorry bout that

OpenStudy (espex):

Yes, @matlee , I dropped a sign, thank you.

OpenStudy (espex):

Your slope is \(-\frac{1}{2}\) sorry for that confusion

OpenStudy (anonymous):

oh so down 1 over 2

OpenStudy (espex):

@ro561man do you recall the equation we were working with previously?

OpenStudy (anonymous):

yeah

OpenStudy (espex):

\(y-y_1=m(x-x_1)\) So what is the 'm', \(x_1\), and \(y_1\)?

OpenStudy (matlee):

No sorry i did not mean to correct you i just had to guess on that one

OpenStudy (anonymous):

m=2 x1=6 y1=3

OpenStudy (espex):

@matlee it's always good to have a second or third pair of eyes when teaching. :) Thank you.

OpenStudy (espex):

@ro561man So the slope for a perpendicular line is....?

OpenStudy (espex):

And put them together in the equation.

OpenStudy (anonymous):

-2

OpenStudy (espex):

Reverse the sign and flip

OpenStudy (anonymous):

oh srry y-3=-2[3-x]

OpenStudy (espex):

\(y-3=-\frac{1}{2}(x-6)\)

OpenStudy (anonymous):

oh it has to b a fraction

OpenStudy (espex):

In some cases, yes. If you have a slope of 3, and you want a perpendicular line, the slope will be \(-\frac{1}{3}\), however, if you have a slope of \(\frac{1}{2}\) and you want the perpendicular, it would be -2.

OpenStudy (espex):

So no matter what it is, you reverse the sign and flip it.

OpenStudy (espex):

Anyway, you now have \(y-3=-\frac{1}{2}(x-6)\), put it into slope/intercept form and you have the answer.

OpenStudy (anonymous):

thank you so much